OpenStudy (anonymous):

Will someone please help show me the process for solving the following equation? solve the equation in the interval [0, 2pi] 2sintcost+sint-2cost-1=0

OpenStudy (anonymous):

First, find the GCF (sin x(2 cos x +1) -1(2 cos x +1) = 0 Then factor (sin x -1)(2 cos x + 1) = 0 Then, sin x - 1 = 0 Use arcsin to solve for x x = arcsin (1) Repeat that for cos Based on that, (sin x-1)(2cos x + 1) = 0, is true when: x = pi/2+-2pin, 2pi/3+-2pin, 4pi/3+-2pin